OpenStudy (anonymous):

1/i=? -1 1 i i^3 i-1

OpenStudy (anonymous):

-i

hartnn (hartnn):

since i*i=-1 1/i=-i then -1 can again be written as i*i so 1/i=i*i*i=cube of i

OpenStudy (phi):

multiply top and bottom by i

OpenStudy (anonymous):

(1*i)/(i)^2

OpenStudy (phi):

what is i^2 ?

OpenStudy (anonymous):

when i multiply the bottom by i, I get i*i, which is the same thing as i^2

OpenStudy (phi):

It is always good to remember that i= sqrt(-1) and i*i= sqrt(-1)*sqrt(-1)= -1 so i^2 is -1